OpenStudy (sloppycanada):

What is the equation of the ellipse with foci (0, 7), (0, -7) and vertices (0, 8), (0, -8)?

OpenStudy (sloppycanada):

\[a^2 + b^2 = c^2\] \[a^2 - c^2 = -b^2\]

OpenStudy (sloppycanada):

8^2 - 7^2 = 15

OpenStudy (sloppycanada):

\[\frac{ x^2 }{ 15 } + \frac{ y^2 }{ 64 } \]